# Noise-Aware Algorithm for Heterogeneous Differentially Private Federated Learning

### v1

@@ 1 Introduction
In the presence of sensitive information in the train data, FL algorithms must be able to provide rigorous data privacy guarantees against a potentially curious server or any third party (Hitaj et al., 2017; Rigaki & García, 2020; Wang et al., 2019b; Zhu et al., 2019; Geiping et al., 2020). Differential Privacy (Dwork et al., 2006b, a; Dwork, 2011; Dwork & Roth, 2014) has been used in DPFL systems to achieve such formal privacy guarantees. When there is a trusted server in the system, FL with central differential privacy (CDP), which is operated by the server by adding controlled noise to the aggregation of clients’ updates, is a solution (McMahan et al., 2018; Geyer et al., 2017). When there is no trusted server, which is more common, FL with local differential privacy (LDP), where each client randomizes its updates locally is also a solution (Zhao et al., 2021). However, LDP is limited in the sense that achieving privacy while preserving model utility is challenging, due to clients’ independent noise additions. Some solutions have been proposed for improving utility in LDP, e.g., using a trusted shuffler system (Liu et al., 2021b; Girgis et al., 2021), which may be difficult to establish if the server itself is not trusted.
Clients often have heterogeneous privacy preferences coming from their varying privacy policies. Furthermore, dataset size usually varies a lot across clients. Additionally, depending on their computational budgets, some clients may use relatively smaller batch sizes locally for running DPSGD algorithm (Abadi et al., 2016). As we will show, a small privacy parameter ($\epsilon$) and/or a small batch size lead to a fast increment of the noise level in a client’s model update. Existing heterogeneous DPFL works mostly either depend on a trusted server, i.e., CDP (Chathoth et al., 2022; Zhou et al., 2022), or suffer from suboptimal and vulnerable aggregation strategies on an untrusted server (i.e., LDP) based on clients’ privacy parameters (Liu et al., 2021a). We consider local heterogeneous DPFL systems with an untrusted server and propose an efficient algorithm, which is aware of the noise level in each client’s model update. We propose to employ Robust PCA (RPCA) algorithm (Candes et al., 2009) by the untrusted server to estimate the amount of noise in clients’ model updates, which we show depends strongly on multiple factors (e.g., their privacy parameter and their batch size ratio), and assign their aggregation weights accordingly. The use of this efficient strategy on the server, which is independent of clients sending any privacy parameters to the server or not, improves model utility and convergence speed while being robust to potential falsifying clients. The highlights of our contributions are the followings:
@@
  • We propose “Robust-HDP”, a noise-aware robust algorithm for local heterogeneous DPFL (untrusted server)
@@ 2 Related work
Figure 1: Security model in local heterogeneous DPFL, where client $i$ has local train data $\mathcal{D}_{i}$ and privacy parameters $(\epsilon_{i},\delta_{i})$, and does not trust any external parties.
@@
The output of an algorithm $\mathcal{M}$, in the sense of Definition 2.1, is all the information that the untrusted server, which we want to protect against, observes. We consider heterogeneous local model of DPFL (Figure 1), where each client $i$ has its own desired privacy parameters ($\epsilon_{i},\delta_{i}$), and sends data-dependent computation results $\mathcal{M}(\mathcal{D}_{i})$ (i.e., model updates) to the server. Also, in the context of Definition 2.1, the notion of neighboring datasets that we consider in this work, refers to pair of federated datasets $d={\{\mathcal{D}_{1},\cdots,\mathcal{D}_{n}\}}$ and $d^{\prime}={\{\mathcal{D}_{1},\cdots,\mathcal{D}_{n}\}}$, differing by one data point of one client (i.e., record-level DPFL).
@@
Liu et al. (2021a) adapted a projection-based approach, similar to that of projected DPSGD (Zhou et al., 2021), to the heterogeneous DPFL setting to propose PFA and improve utility. Although assuming an untrusted server, their proposed algorithm relies on the assumption that the server knows the clients’ “true” privacy parameters $\{(\epsilon_{i},\delta_{i})\}$ and uses them to cluster clients to “public” (those with larger privacy parameters) and “private”. As such, as we show, PFA is extremely vulnerable to when clients share a falsified value of their privacy parameters (often larger than their true values) with server. Also, they used aggregation strategy $w_{i}\propto\epsilon_{i}$ on server for PFA and another algorithm called WeiAvg (see Table 1). As we will show, even if the server knows clients’ true privacy parameters, this information is not a perfect indication of the “true” noise level in their model updates, especially with heterogeneous privacy parameters and batch/dataset sizes.
The current state of the art in local heterogeneous DPFL calls for a robust algorithm that takes all the mentioned potential sources of heterogeneity across clients into account and achieves high utility and data privacy simultaneously.
@@ 3.4 Description of Robust-HDP algorithm
5             $w_{i}^{e}\leftarrow\frac{{1}/{\|\mathbf{S}_{:,i}\|_{2}^{2}}}{\sum_{j\in\mathcal{S}_{e}}{1}/{\|\mathbf{S}_{:,j}\|_{2}^{2}}}$
@@ H.2 Local DPFL with highly heterogeneous data split across clients (future work)
On the other hand, we have DPFL systems with highly heterogeneous data splits. In the existence of a trusted server (CDP), an idea was proposed by Chathoth et al. (2022) for clustering clients with cohort-level privacy with privacy and data heterogeneity across cohorts, using $\epsilon$-DP definition (Definition 2.1 with $\delta=0$). When there is no trusted server (LDP), we can follow a similar direction of clustered DPFL to address scenarios with highly heterogeneous data splits: clients are first clustered by the server such that the data distribution of clients in a cluster are more similar to each other, and then, a model is learned for each cluster. However, the DP noise in clients’ model updates makes clustering of clients harder. A recent work in (Malekmohammadi et al., 2024) has addressed this scenario by proposing an algorithm, which is robust to the DP noise existing in clients’ model updates, for clustering clients in DPFL system with local differential privacy.

### v3

@@ 1 Introduction
In the presence of sensitive information in the train data, FL algorithms must be able to provide rigorous data privacy guarantees against a potentially curious server or any third party (Hitaj et al., 2017; Rigaki & García, 2020; Wang et al., 2019b; Zhu et al., 2019; Geiping et al., 2020). Differential Privacy (Dwork et al., 2006b, a; Dwork, 2011; Dwork & Roth, 2014) has been used in DPFL systems to achieve such formal privacy guarantees. When there is a trusted server in the system, DP is provided by the server adding controlled noise to the aggregation of clients’ updates (McMahan et al., 2018; Geyer et al., 2017). When there is no trusted server, local perturbations, where each client randomizes its updates locally before sending them to the server, is also a solution (Zhao et al., 2021). However, this method is limited in the sense that achieving privacy while preserving model utility is challenging, due to clients’ independent local noise additions. Some solutions have been proposed for improving utility of this method, e.g., using a trusted shuffler system (Liu et al., 2021b; Girgis et al., 2021), which may be difficult to establish if the server itself is not trusted.
Clients often have heterogeneous privacy preferences coming from their varying privacy policies. Furthermore, dataset size usually varies a lot across clients. Additionally, depending on their computational budgets, some clients may use relatively smaller batch sizes locally for running DPSGD algorithm (Abadi et al., 2016). As we will show, a small privacy parameter ($\epsilon$) and/or a small batch size lead to a fast increment of the noise level in a client’s model update. Existing heterogeneous DPFL works mostly either depend on a trusted server (Chathoth et al., 2022; Zhou et al., 2022), or suffer from suboptimal and vulnerable aggregation strategies on an untrusted server (based on clients’ privacy parameters (Liu et al., 2021a)). We consider a heterogeneous DPFL systems with an untrusted server and propose an efficient aggregation strategy for the server, which is aware of the noise level in each client’s model update: we propose to employ Robust PCA (RPCA) algorithm (Candes et al., 2009) by the untrusted server to estimate the amount of noise in clients’ model updates, which we show depends strongly on multiple factors (e.g., their privacy parameter and their batch size ratio), and assign their aggregation weights accordingly. The use of this efficient strategy on the server, which is independent of clients sending any privacy parameters to the server or not, improves model utility and convergence speed while being robust to potential falsifying clients. The highlights of our contributions are the followings:
@@
  • We propose “Robust-HDP”, a noise-aware robust algorithm for heterogeneous DPFL.
@@ 2 Related work
Figure 1: Security model in heterogeneous DPFL, where client $i$ has local train data $\mathcal{D}_{i}$ and privacy parameters $(\epsilon_{i},\delta_{i})$, and does not trust any external parties.
@@
The output of an algorithm $\mathcal{M}$, in the sense of Definition 2.1, is all the information that the untrusted server, which we want to protect against, observes. We consider heterogeneous DPFL (Figure 1), where each client $i$ has its own privacy parameters ($\epsilon_{i},\delta_{i}$), and sends data-dependent computation results $\mathcal{M}(\mathcal{D}_{i})$ (model updates) to the server. Also, in the context of Definition 2.1, the notion of neighboring datasets that we consider in this work, refers to pair of federated datasets $d={\{\mathcal{D}_{1},\cdots,\mathcal{D}_{n}\}}$ and $d^{\prime}={\{\mathcal{D}_{1},\cdots,\mathcal{D}_{n}\}}$, differing by one data point of one client (i.e., record-level DPFL).
@@
Liu et al. (2021a) adapted a projection-based approach, similar to that of projected DPSGD (Zhou et al., 2021), to the heterogeneous DPFL setting to propose PFA and improve utility. Although assuming an untrusted server, their proposed algorithm relies on the assumption that the server knows the clients’ “true” privacy parameters $\{(\epsilon_{i},\delta_{i})\}$ and uses them to cluster clients to “public” (those with larger privacy parameters) and “private”. As such, as we show, PFA is extremely vulnerable to when clients share a falsified value of their privacy parameters (often larger than their true values) with server. Also, they used aggregation strategy $w_{i}\propto\epsilon_{i}$ on server for PFA and another algorithm called WeiAvg (see Table 1). As we will show, even if the server knows clients’ true privacy parameters, this information is not a perfect indication of the “true noise level” in their model updates, especially with heterogeneous privacy parameters and batch/dataset sizes.
The current state of the art in heterogeneous DPFL calls for a robust algorithm that takes all the mentioned potential sources of heterogeneity across clients into account and achieves high utility and data privacy simultaneously.
@@ 3.4 Description of Robust-HDP algorithm
5             $w_{i}^{e}\leftarrow\frac{{1}/{\|\mathbf{S}_{:,i}\|_{2}^{2}}}{\sum_{j\in\mathcal{S}^{e}}{1}/{\|\mathbf{S}_{:,j}\|_{2}^{2}}}$
6      $\bm{\theta}^{e+1}\leftarrow\bm{\theta}^{e}+\sum_{i\in\mathcal{S}^{e}}w_{i}^{e}\Delta\tilde{\bm{\theta}}_{i}^{e}$
@@ H.2 Local DPFL with highly heterogeneous data split across clients (future work)
On the other hand, we have DPFL systems with highly heterogeneous data splits. In the existence of a trusted server, an idea was proposed by Chathoth et al. (2022) for clustering clients with cohort-level privacy with privacy and data heterogeneity across cohorts, using $\epsilon$-DP definition (Definition 2.1 with $\delta=0$). When there is no trusted server, we can follow a similar direction of clustered DPFL to address scenarios with highly heterogeneous data splits: clients are first clustered by the server such that the data distribution of clients in a cluster are more similar to each other, and then, a model is learned for each cluster. However, the DP noise in clients’ model updates makes clustering of clients harder. A recent work in (Malekmohammadi et al., 2024) has addressed this scenario by proposing an algorithm, which is robust to the DP noise existing in clients’ model updates, for clustering clients in DPFL system with local differential privacy.